need to solve system by elimination.
7x - y = 14
6x + y = 25

Answer by checkley79(3341)

7x - y = 14
6x + y = 25 add.
-----------------------
13x=39x=39/3
x=3 ans.
7*3-y=14
21-y=24
-y=14-21
-y=-7
y=7 ans.
Proof:
6*3*7=25+
18+7=25
25=25